OpenStudy (anonymous):

Solve equation 10b + 18 = 8b + 4

OpenStudy (ajprincess):

Subtract 18 both sides. what do u get if u do so @mikee118?

OpenStudy (anonymous):

Idk

OpenStudy (whpalmer4):

\[10b + 18 - 18 = 8b + 4 - 18\]\[10b = 8b -14\]Now try the same thing with 8b.

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

Need help

OpenStudy (anonymous):

Idk

OpenStudy (whpalmer4):

Okay, you've got b on both sides of the equation. You want it on one side, by itself. You can move things around by adding them to both sides of the equation, or subtracting them from both sides. You saw how we moved the 18 from the left side to the right side by subtracting it from both sides. Now we are going to subtract 8b from both sides. Try it.

OpenStudy (aravindg):

@mikee118 did you get it now?

OpenStudy (anonymous):

10b+18=8b+4 so in this equation you have to move similar values on the same sides therefore moving 18 to the right making it negative 18 and move 8b to the left which turns out as negative 8b. Now you get this equation: 10b-8b=4-18 >so now you continue to solve.. 2b=-14 >divide the equation by 2 you get.. b=-7

OpenStudy (aravindg):

@cozyfire We at openstudy do not appreciate giving out worked out direct answer like that ..instead help the user get to the answer like others were trying above !That will help the asker more ....worth reading Openstudy COC http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ahh i see so this is a kind of tutoring forum or something i understand sorry i won't do it again..i've been in this site for 2 days so i don't really know much

OpenStudy (aravindg):

yep i understand you are a newbie here :) ..next time ensure you make the asker also participate in the question....Thanks for understanding :)